OpenStudy (anonymous):

check my answer please will medal a herd of zebra currently has 45 members. based on the available resources biologists estimate that the size of the herd will increase at a rate of 7% per year. which of the following graphs models this relationship if the x axis represents years and the y axis represents number of zebra.

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

so when it says it increases at a rate of 7% per year, what kind of function is that?

OpenStudy (danjs):

Hint: It is section 3 in your review list on that web site.

OpenStudy (danjs):

THe initial condition is 45 zebras. when x=0, y=45

OpenStudy (danjs):

What will y(1) = The initial amount increases by 7% each increase of 1 unit in x years

OpenStudy (anonymous):

exponential

OpenStudy (danjs):

right

OpenStudy (anonymous):

48.15

OpenStudy (danjs):

Remember the general form looks something like y(x) = A*b^x

OpenStudy (danjs):

right, 48.15 is y(1), the number of zebras after 1 year

OpenStudy (danjs):

Here is how to form the equation to use....

OpenStudy (danjs):

\[y(x) = A*b^x\] They tell you that initially there are 45 zebras. That is when x=0. The point is x=0, y=45

OpenStudy (danjs):

\[y(0) = 45 = A*b^0\]

OpenStudy (danjs):

look familiar?

OpenStudy (anonymous):

yeah

OpenStudy (danjs):

a base to the zero power equals 1, so that point x=0, y=45 gives you the value for A y(0) = 45 = A *1 A =45

OpenStudy (danjs):

\[y(x) = 45*b^x\] Now just have to find b.

OpenStudy (danjs):

you could use another data point. You figured y(1) = 48.15 zebras.

OpenStudy (danjs):

y(1) = 48.15 = 45(b)^1 b = 1.07

OpenStudy (danjs):

Overall function is then \[y(x) = 45*1.07^x\]

OpenStudy (danjs):

To pick what graph, let x be a higher number of years, x = 20, what is the population at 20 years? y(20) = 45*1.07^20 = 174.13 zebras

OpenStudy (danjs):

Graph B has the axis labeled wrong You chose the right one, D

OpenStudy (danjs):

@Dianie

OpenStudy (anonymous):

ok so d was correct thanks